OpenStudy (nudeoua):

Below are the steps to solve an equation: Step 1: |x − 5| + 2 = 5 Step 2: |x − 5| = 5 − 2 Step 3: |x − 5| = 3 Which of the following is a correct next step to solve the equation? x + 5 = −3 −x − 5 = 3 x − 5 = −3 x + 5 = 3 I think its the second one.

imqwerty (imqwerty):

after the 3rd step we open the mod value when we open the mod value it gives us 2 equtions- |x-5|=3 +(x-5)=3 and -(x-5)=3 x+5=3 5-x=3